OpenStudy (anonymous):

14. Write an equation for the line through the point P(4, 7) with a slope of -3 in slope-intercept form.

OpenStudy (nurali):

An equation of a line through a point P(x1 , y1) with slope m is given by y-y1=m(x-x1) Point (4,7), then x1=4 and y1=7 The slope is m=-3 y-7=-3(x-(4)) y-7=-3(x-4) y=-3x+12+7 y=-3x-19 the equation of a line

OpenStudy (anonymous):

which part of that above is the answer @nurali